OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 4x + y = 11 x + 2y = 8

OpenStudy (anonymous):

x = 8-2y. Substitute this in the first equation to get an equation in x to solve.

OpenStudy (anonymous):

x=8-2y 4(8-2y)+y=11 32-8y+y=11 21=7y y=3 x=8-2y x=8-2(3) x=2